Which of the following compounds will have the highest molar solubility in pure water?

  1. PbSO4, Ksp = 1.82
Answer #1

solubility\; \; \; = \sqrt{k_{sp}}

so sulubility will be higher for that salt which will have higher ksp

and MgCO3 has higher ksp among the all of the above options
Answer #2

1)Ksp of PbSO4 = (solubility)^2

or, solubility = 1.35*10^-4 M

2) Ksp of MgCO3 = (solubility)^2

or, solubility = 2.612*10^-3 M

3) Ksp of AgI = (solubility)^2

or, solubility = 9.225*10^-9 M

4) Ksp of PbS = (solubility)^2

or, solubility = 9.51*10^-15 M

5) Ksp of FeS = (solubility)^2

or, solubility = 6.12*10^-10 M

thus, MgCO3 has the highest solubility